A body P is thrown vertically upwards . Another body Q of same mass thrown at an angle 60 . If both attain same maximum height , then ratio of initial kinetic energy of P to Q is what ?

Grade:11

1 Answers

Prsahant Parashar
7 years ago
h=v2sin x/2g
 
for (1) sin x = 90  and (2) sin x=60
now 1
h=v21/2g   {sin 90=1}...........1
 
2.
h=3v22/8g.......2 {sin 60= 31/2/2}
 
    from..   1/2
we get v21/v22=3/4
this is the ratio of kinttic energy since mass is same